Given logb3=.72 and logb5=1.05
Find logb.6

Answer by stanbon(75887)

Given logb3=.72 and logb5=1.05
Find logb.6
-----
logb(0.6) = logb(3/5) = logb(3)-logb(5) = 0.72-1.05 = -0.33
------------
Cheers,
Stan H.
------------